# RIZE: Regularized Imitation Learning via Distributional Reinforcement Learning

### v2

@@ 4.1 Distributional Value Integration
Our approach departs from traditional imitation learning by explicitly using value distributions as critics within an actor-critic framework (Zhou et al., 2023). We argue that learning the soft value distribution $Z(s,a)$ in Equation (1), rather than relying solely on point estimates like $Q(s,a)$, enhances decision-making by capturing uncertainty in complex environments. This view is consistent with recent neuroscience findings suggesting that decision-making in the prefrontal cortex relies on learning distributions over outcomes rather than only their expectations (Muller et al., 2024).
Moreover, access to the full return distribution enables the computation of statistical moments—most notably the expectation—which we optimize both policy and critic using the expectation of $Z$ (Bellemare et al., 2017; Dabney et al., 2018b; a), yielding a more robust learning signal while remaining compatible with IQ-Learn.
We compute $Q$ as the expectation of the soft value distribution:
@@ 4.2 Implicit Reward Regularization
In this section, we propose a regularizer for inverse RL that refines existing implicit reward formulations (Garg et al., 2021). The implicit reward is defined as:
@@
Previous works typically regularize implicit rewards either using $L_{2}$-norms (Garg et al., 2021) or by treating them as squared-TD errors between rewards and fixed targets (Reddy et al., 2020; Al-Hafez et al., 2023). While we adopt a similar squared-TD setting, we introduce adaptive targets $\lambda^{\pi_{E}}$ (for the expert $\pi_{E}$) and $\lambda^{\pi}$ (for the imitation policy $\pi$) to construct our convex regularizer $\Gamma:\mathbb{R}^{S\times A}\rightarrow\bar{\mathbb{R}}$:
@@
As in IQ-Learn and LSIQ, our method seeks behavior indistinguishable from expert demonstrations in a non-adversarial implicit-reward setting. Prior work analyzes optimal implicit rewards, drawing on Max–Min analyses from GANs (Al-Hafez et al., 2023; Goodfellow et al., 2014). Because we bind rewards to adaptive targets via $\Gamma(R_{Q},\lambda)$ (Equation (7)), we carry out an analogous analysis, stated below.
@@
As established in Corollary A.2, this optimal reward satisfies:
@@
where $\lambda_{\min}:=\min\left\{\lambda^{\pi_{E}},\lambda^{\pi}\right\}$ and $\lambda_{\max}:=\max\left\{\lambda^{\pi_{E}},\lambda^{\pi}\right\}$. The coefficient $c$ and targets bound rewards within a well-defined range. In practice, we use $\lambda\in[0,10]$ and $c\in[0.1,0.5]$, which we found to promote stable training (see Appendix B.3).
@@
Recalling that we represent rewards through Q-values, the boundedness of the optimal reward ensures that critic updates remain constrained and—because the policy directly depends on these critic values—promotes stable policy optimization. However, as previously noted (Al-Hafez et al., 2023; Viano et al., 2022), the convergence guarantee originally stated for IQ-Learn (Garg et al., 2021) does not extend to the $\chi^{2}$-regularizer used in practice. Consequently, a formal proof for the resulting alternating SAC updates remains an open question and is left for future work.
@@ 4.3 Practical Algorithm
We now introduce RIZE: Regularized Imitation Learning via Distributional Reinforcement Learning (Algorithm 1). We approximate $Z$ and $\pi$ with neural networks, optimizing via $Q=\mathbb{E}[Z]$. Following Distributional SAC (Ma et al., 2020), we employ target policies for next-action sampling and a double-critic architecture with target networks for stability. Lower learning rates proved critical for robust policy updates, with a four-layer MLP policy network necessary for complex tasks.
@@ 5 Experiments
We evaluate our algorithm on four MuJoCo (Todorov et al., 2012) benchmarks (HalfCheetah-v2, Walker2d-v2, Ant-v2, Humanoid-v2) against state-of-the-art imitation learning methods: IQ-Learn (Garg et al., 2021), LSIQ (Al-Hafez et al., 2023), and SQIL (Reddy et al., 2020). All experiments are conducted with five seeds for statistical significance (Henderson et al., 2018). We assess each method using three and ten expert trajectories. For ten trajectories, we retain baseline hyperparameters; for three trajectories (not reported in prior works), we adapt configurations from single-demonstration settings. IQ-Learn and LSIQ use their official implementations, while SQIL is evaluated using the LSIQ codebase. Results report mean performance across five seeds, with half a standard deviation to indicate uncertainty, and lines are smoothed for better readability. We normalize episode returns based on expert performance. For details, expanded visualizations, and hyperparameter tuning, see Appendix B.
Figure 1: Normalized return of RIZE versus online imitation learning baselines on Gym MuJoCo tasks. We depict the sorted top 25% episodic returns across five seeds to evaluate convergence to expert-level performance, using three and ten expert trajectories.
5.1 Main Results
Our method outperforms LSIQ and SQIL across tasks, with IQ-Learn as the only competitive baseline. Notably, in the most complex environment, Humanoid-v2, our approach is the sole method achieving expert-level performance, while all baselines fail (see Figures 1, 2). This demonstrates our algorithm’s scalability to high-dimensional control problems. Additionally, our method shows superior sample efficiency, requiring fewer gradient steps to match expert performance compared to SOTA algorithms. These results highlight the robustness and efficiency of our approach in tackling complex tasks.
Figure 2: Normalized return of RIZE vs. online imitation learning baselines on Gym MuJoCo tasks. n denotes the number of expert trajectories.
5.2 Recovered Reward
Here, we compare the reward behavior of RIZE against the baselines. As shown in Figure 3, our approach effectively bounds the rewards for both expert and policy samples across environments. The reward regularizer $\Gamma(R_{Q},\lambda)$ keeps the recovered rewards close to adaptive targets, with the optimization of $\lambda^{\pi_{E}}$ and $\lambda^{\pi}$ via Equation (8) enforcing an upper bound on expert rewards and a lower bound on policy rewards. With sufficient expert trajectories, rewards for both expert and policy samples stabilize around consistent values. In contrast, baseline methods exhibit unbounded reward growth in the Humanoid task, and SQIL fails to constrain rewards even in the simpler HalfCheetah environment.
Figure 3: Implicit reward curves for expert and policy samples on MuJoCo tasks. We evaluate on ten expert demonstrations.
5.3 Role of Distributional Value Function
We examine the evolution of critics in RIZE compared to the baselines. Unlike the baselines, which train Q-networks, RIZE employs distributional value functions and uses their expectation $\mathbb{E}[Z(s,a)]$ for loss optimization. Since all methods update policies by maximizing state-action values, unbounded values can significantly undermine robustness. As illustrated in Figure 4, RIZE and LSIQ maintain bounded state-action values in the challenging Humanoid environment, whereas SQIL and IQ-Learn do not. Nevertheless, the boundedness of critics alone does not guarantee expert-like policy performance, as evidenced by LSIQ’s results in Figure 2.
Figure 4: State-action value curves for expert and policy samples on MuJoCo tasks. Ten trajectories are used throughout the analysis.
@@ 5.3 Role of Distributional Value Function
Moreover, to assess the significance of quantile-based value networks Z(s,a)\,Z(s,a), we train RIZE with point-estimate critics Q(s,a)\,Q(s,a) and compare the results with the original implementation. As shown in Figure 5, distributional value functions consistently outperform Q-networks across all tasks, exhibiting lower variance and greater sample efficiency. Notably, in the Humanoid environment, Q-values fail to replicate expert demonstrations, whereas quantile-based critics maintain expert-level performance.
Figure 5: Normalized return of RIZE on Gym MuJoCo tasks utilizing value distribution $Z(s,a)$ or classic $Q(s,a)$ as the critic. We use three expert trajectories for all tasks.
5.4 Ablation: Regularization Strategies
We investigate the effect of squared-TD error regularization on the Ant-v2 task using three distinct formulations: Expert-focused TD: $\mathbb{E}_{\rho_{E}}[R_{Q}-\lambda^{\pi_{E}}]^{2}+\mathbb{E}_{\rho_{\pi}}[R_{Q}]^{2}$, Policy-focused TD: $\mathbb{E}_{\rho_{E}}[R_{Q}]^{2}+\mathbb{E}_{\rho_{\pi}}[R_{Q}-\lambda^{\pi}]^{2}$, Baseline L2: $\mathbb{E}_{\rho_{E}}[R_{Q}]^{2}+\mathbb{E}_{\rho_{\pi}}[R_{Q}]^{2}$.
Our results in Figure 6 demonstrate that applying squared-TD regularization to both expert and policy samples improves robustness. While the expert-focused TD formulation outperforms the policy-focused variant, neither consistently achieves expert-level behavior. Over time, both exhibit gradual performance degradation, underscoring the challenges of maintaining stable imitation. This highlights the need for more refined regularization strategies to ensure consistent and reliable learning in imitation learning.
Figure 6: Ablation study on the effect of various regularization strategies, evaluated using three demonstrations.
@@ 6 Conclusion
We propose a novel IRL framework that overcomes the limitations of fixed reward mechanisms through dynamic reward adaptation and context-sensitive regularization. Our approach ensures bounded implicit rewards and stable value function updates, leading to robust policy optimization. By integrating distributional RL with implicit reward learning, we capture richer return dynamics while preserving theoretical guarantees. Empirical results on MuJoCo benchmarks show expert-like proficiency on the Humanoid task with only three demonstrations. Ablation studies confirm the critical role of our regularization mechanism. This work unifies implicit reward regularization with distributional return representations, offering a scalable and sample-efficient solution for complex decision-making. Future directions include extending this framework to offline IL and risk-sensitive robotics, where adaptive rewards and uncertainty-aware distributions can further improve robustness and generalization.
@@ Appendix A Supporting Proofs
A.1 Expectation of the Value Distribution
@@
From the soft value distribution definition
@@
A.2 Bounds on the Optimal Implicit Reward
Corollary A.2.
The optimal implicit reward satisfies:
where $\lambda_{\min}:=\min\left\{\lambda^{\pi_{E}},\lambda^{\pi}\right\}$ and $\lambda_{\max}:=\max\left\{\lambda^{\pi_{E}},\lambda^{\pi}\right\}$.
Proof.
From Proposition 4.1:
$R_{Q}^{*}(s,a)=\underbrace{\frac{\rho_{E}-\rho_{\pi}}{2c(\rho_{E}+\rho_{\pi})}}_{\text{Term 1}}+\underbrace{\frac{\rho_{E}\lambda^{\pi_{E}}+\rho_{\pi}\lambda^{\pi}}{\rho_{E}+\rho_{\pi}}}_{\text{Term 2}}\,.$ (16)
Term 1 lies in $\left[-\frac{1}{2c},\;\frac{1}{2c}\right]$ since $\rho_{E},\rho_{\pi}\geq 0$ (achieved when either $\rho_{\pi}\to 0$ or $\rho_{E}\to 0$). Term 2 is a convex combination of $\lambda^{\pi_{E}}$ and $\lambda^{\pi}$, thus in $\left[\lambda_{\min},\;\lambda_{\max}\right]$. Combining intervals yields the result. ∎
A.3 Convergence of the Optimal Reward
@@
Appendix B Additional Experiments
B.1 MuJoCo Control Suite
@@
We evaluate our imitation learning approach, RIZE, on four benchmark Gym (Brockman et al., 2016) MuJoCo (Todorov et al., 2012) environments: HalfCheetah-v2, Walker2d-v2, Ant-v2, and Humanoid-v2. Expert trajectories were obtained from IQ-Learn (Garg et al., 2021), generated by a Soft Actor-Critic agent (Haarnoja et al., 2018), with each trajectory comprising 1,000 state-action transitions. To facilitate performance comparisons, episode returns are normalized relative to expert performance: HalfCheetah (5100), Walker2d (5200), Ant (4700), and Humanoid (5300). All implementation details used in our experiments are publicly available at https://github.com/adibka/RIZE.
B.2 Implementation Details
@@
Our architecture integrates components from Distributional SAC (DSAC) [^2] https://github.com/xtma/dsac (Ma et al., 2020) and IQ-Learn [^3] https://github.com/Div99/IQ-Learn (Garg et al., 2021), with hyperparameters tuned through systematic search and ablation studies. Key configurations for experiments involving three and ten demonstrations are summarized in Table 1.
@@
Distributional SAC Components: The critic network is implemented as a three-layer multilayer perceptron (MLP) with 256 units per layer, trained using a learning rate of $3\times 10^{-4}$. The policy network is a four-layer MLP, also with 256 units per layer. Environment-specific learning rates are applied: $1\times 10^{-5}$ for the Humanoid environment and $5\times 10^{-5}$ for all others. To enhance training stability, we employ a target policy—a delayed version of the online policy—and sample next-state actions from this module. For value distribution training $Z^{\pi}_{\phi,\tau}$, we adopt the Implicit Quantile Networks (IQN) (Dabney et al., 2018a) approach by sampling quantile fractions $\tau$ uniformly from $\mathcal{U}(0,1)$. Additionally, dual critic networks with delayed updates are used, which empirically improve training stability.
@@
IQ-Learn Adaptations: Key adaptations from IQ-Learn include adjustments to the regularizer coefficient $c$ and entropy coefficient $\alpha$. Specifically, for the regularizer coefficient $c$, we find that $c=0.5$ yields robust performance on the Humanoid task, while $c=0.1$ works better for other tasks. For the entropy coefficient $\alpha$, smaller values lead to more stable training. Unlike RL, where exploration is crucial, imitation learning relies less on entropy due to the availability of expert data. Across all tasks, we set initial target reward parameters as $\lambda^{\pi_{E}}=10$ and $\lambda^{\pi}=5$. Furthermore, we observe that lower learning rates for target rewards improve overall learning performance.
@@
The choice between v0 or value loss variants depends on environment complexity: we find that for a complex task like Humanoid-v2, the v0 variant demonstrates greater robustness. And, for HalfCheetah-v2, Walker2d-v2, and Ant-v2, the value variant performs better.

### v3

@@ 4.1 Distributional Value Integration
Our approach departs from traditional imitation learning by explicitly using return distributions as critics within an actor-critic framework (Zhou2023). We argue that learning the soft return distribution $Z(s,a)$ in Equation (1), rather than relying solely on point estimates like $Q(s,a)$, enhances decision-making by capturing uncertainty in complex environments. This view is consistent with recent neuroscience findings suggesting that decision-making in the prefrontal cortex relies on learning distributions over outcomes rather than only their expectations (Muller2024).
Moreover, access to the full return distribution enables the computation of statistical moments—most notably the expectation—which we optimize both policy and critic using the expectation of $Z$ (Bellemare2017; Dabney2018a; Dabney2018), yielding a more robust learning signal while remaining compatible with IQ-Learn.
We compute $Q$ as the expectation of the soft return distribution:
@@ 4.2 Implicit Reward Regularization
In this section, we propose a regularizer for inverse RL that refines existing implicit reward formulations (Garg2021). The implicit reward is defined as:
@@
Previous works typically regularize implicit rewards either using $L_{2}$-norms (Garg2021) or by treating them as squared-TD errors between rewards and fixed targets (Reddy2020; Al-Hafez2023). While we adopt a similar squared-TD setting, we introduce adaptive targets $\lambda^{\pi_{E}}$ (for the expert $\pi_{E}$) and $\lambda^{\pi}$ (for the imitation policy $\pi$) to construct our convex regularizer $\Gamma:\mathbb{R}^{S\times A}\rightarrow\bar{\mathbb{R}}$:
@@
As in IQ-Learn and LSIQ, our method seeks behavior indistinguishable from expert demonstrations in a non-adversarial implicit-reward setting. Prior work analyzes optimal implicit rewards, drawing on Max–Min analyses from GANs (Al-Hafez2023; Goodfellow2014). Because we bind rewards to adaptive targets via $\Gamma(R_{Q},\lambda)$ (Equation (7)), we carry out an analogous analysis, stated below.
@@
Corollary 4.2.
The optimal implicit reward satisfies:
@@
where $\lambda_{\min}:=\min\left\{\lambda^{\pi_{E}},\lambda^{\pi}\right\}$ and $\lambda_{\max}:=\max\left\{\lambda^{\pi_{E}},\lambda^{\pi}\right\}$. The coefficient $c$ and adaptive targets bound rewards within a well-defined range.
Proof.
Considering the optimal reward Equation (10), the first term lies in $\left[-\frac{1}{2c},\;\frac{1}{2c}\right]$ since $\rho_{E},\rho_{\pi}\geq 0$ (achieved when either $\rho_{\pi}\to 0$ or $\rho_{E}\to 0$). The second term is a convex combination of $\lambda^{\pi_{E}}$ and $\lambda^{\pi}$, thus in $\left[\lambda_{\min},\;\lambda_{\max}\right]$. Combining intervals yields the result. In practice, we use $\lambda\in[5,10]$ and $c\in[0.1,0.5]$, which we found to promote stable training (see Appendix C.7). ∎
@@
Recalling that we represent rewards through Q-values, the boundedness of the optimal reward ensures that critic updates remain constrained and—because the policy directly depends on these critic values—promotes stable policy optimization. However, as previously noted (Al-Hafez2023; Viano2022), the convergence guarantee originally stated for IQ-Learn (Garg2021) does not extend to the $\chi^{2}$-regularizer used in practice. Consequently, a formal proof for the resulting alternating SAC updates remains an open question and is left for future work.
@@ 4.3 Practical Algorithm
We now introduce RIZE: Adaptive Regularization for Imitation Learning (Algorithm 1). We approximate $Z$ and $\pi$ with neural networks, optimizing via $Q=\mathbb{E}[Z]$. Following Distributional SAC (Ma2020), we employ target policies for next-action sampling and a double-critic architecture with target networks for stability. Lower learning rates proved critical for robust policy updates, with a four-layer MLP policy network necessary for complex tasks.
@@
Demo 3
Demo 10
Figure 1: RLiable (agarwal2021deep) plots for RIZE vs. BC, LSIQ, SQIL, CSIL, and IQ-Learn on six MuJoCo/Adroit tasks. For each setting (3 demos; 10 demos), we report aggregate Median, IQM, Mean, and Optimality Gap with 95% confidence intervals computed via percentile bootstrap stratified over tasks and five seeds. Scores are normalized to expert performance. Higher is better for Median, IQM, and Mean; lower is better for Optimality Gap.
@@ 5 Experiments
We study continuous-control imitation learning from state–action expert samples, evaluating our algorithm on five MuJoCo (Todorov2012) benchmarks (HalfCheetah-v2, Walker2d-v2, Ant-v2, Humanoid-v2, Hopper-v2) and one Adroit Hand task (Hammer-v1). We compare against state-of-the-art baselines IQ-Learn (Garg2021), LSIQ (Al-Hafez2023), SQIL (Reddy2020), CSIL (Watson2023) and Behavior Cloning (BC) (Pomerleau1991). All experiments use five random seeds (Henderson2018). We assess each method with three and ten expert trajectories. We report mean $\pm$ 95% confidence intervals (CI) across five seeds. Episode returns are normalized by expert performance. For implementation details, additional experimental results, and visualizations, see Appendix B and C.
Figure 2: Normalized returns on MuJoCo and Adroit tasks for RIZE and baselines. We first compute, per seed, the average episodic return over the final third of training steps; bars show the mean across five seeds and error bars denote the 95% confidence interval. Returns are normalized to expert performance and reported for both 3 and 10 expert demonstrations.
Main Results.
Figure 1 reports aggregate metrics computed with RLiable (agarwal2021deep), which provides statistically principled evaluation via Median, Interquartile Mean (IQM), Mean, and Optimality Gap with stratified bootstrap confidence intervals; we adopt this protocol throughout, consistent with recent IL practice (e.g., AILBoost (chang2024adversarial); SFM (jain2025nonadversarial)). Across both 3- and 10-demonstration settings, RIZE achieves higher Median, IQM, and Mean and a lower Optimality Gap than the baselines; the only competitive methods overall are CSIL and IQ-Learn. Increasing the number of expert trajectories from 3 to 10 consistently improves RIZE on all RLiable aggregates. Task-wise, RIZE is the only method that solves Humanoid-v2 (all baselines fail), while on Hammer-v1 RIZE, CSIL, and BC show higher returns than the remaining methods. In terms of sample efficiency, CSIL often attains strong returns early—benefiting from behavior-cloning initialization—but when BC is ineffective (e.g., Humanoid-v2), CSIL underperforms whereas RIZE ultimately succeeds. These trends are corroborated by the bar plot in Figure 2 (average of the final third of training) and the learning curves in Figure 3. For learning curves with three demonstrations, see Figure 6.
Figure 3: Learning curves on MuJoCo and Adroit tasks with 10 expert demonstrations. Lines show the mean normalized return across five seeds; shaded regions denote 95% confidence intervals.
Recovered Rewards.
In this section, we assess whether our choice of regularizer $\Gamma(R_{Q},\lambda)$ in Equation (7) with adaptive targets can effectively bound the implicit rewards. By Corollary 4.2 together with Equation (12), the optimal implicit reward is confined to the interval $\left[-\frac{1}{2c}+\lambda_{\min},\;\frac{1}{2c}+\lambda_{\max}\right]$. Since expert rewards are supposed to be higher than those of the learner during training (based on IRL objective loss 4), and the adaptive targets track these levels, we empirically observe $\lambda^{\pi_{E}}\geq\lambda^{\pi}$; hence $\lambda_{\max}=\lambda^{\pi_{E}}$ and $\lambda_{\min}=\lambda^{\pi}$. In practice, we should see rewards confined to $\left[-\frac{1}{2c}+\lambda^{\pi},\;\frac{1}{2c}+\lambda^{\pi_{E}}\right]$.
Figure 4 depicts the recovered reward trajectories together with the theoretical upper and lower bounds. Across tasks, the curves remain within the predicted band, indicating that our regularizer with adaptive targets effectively bounds both expert and policy rewards. We also observe that the position and width of the band vary with the task and data regime, reflecting the flexibility of the adaptive targets to adjust to different dynamics. In contrast, fixed (non-adaptive) targets lack this ability, leading either to overly loose bounds or to target–reward mismatch. Overall, these results support the role of $\Gamma(R_{Q},\lambda)$ and the adaptive targets in producing stable, well-calibrated implicit rewards that align with Corollary 4.2. For the 3-demonstration setting, Figure 7 shows reward trajectories with their theoretical bounds; for cross-method comparisons of reward trajectories, see Figures 9, 8, and 10(a).
Figure 4: Implicit reward curves for expert and policy samples on MuJoCo and Adroit tasks with 10 expert demonstrations. Each subplot reports the mean across five seeds, with shaded regions showing the 95% confidence interval. Theoretical upper and lower bounds derived in this work are overlaid as separate curves in each subplot.
Ablation on Critic Architecture.
@@
To assess the effect of modeling the return distribution, we replace the IQN-based critic $Z(s,a)$ (Dabney2018) in RIZE with a point-estimate $Q(s,a)$ critic and compare against the original implementation. Here, $Z(s,a)$ denotes the full distribution of discounted returns whose expectation yields $Q(s,a)$. As shown in Figure 5, the IQN critic consistently outperforms the $Q$-network across all tasks, exhibiting lower variance and greater sample efficiency. Notably, on Humanoid-v2 and Hammer-v1, the $Q$-network fails to match expert performance, whereas the IQN critic maintains expert-level returns.
See Figure 13 for reward trajectories with theoretical bounds when using a $Q$-network in place of the IQN critic, and Figures 11, 12, and 10(b) for critic value–estimation curves on MuJoCo and Adroit tasks.
Figure 5: Ablation on critic architecture: $Z(s,a)$ via Implicit Quantile Networks (IQN) (Dabney2018) versus classic $Q(s,a)$. We report expert–normalized returns across all MuJoCo and Adroit tasks using three expert demonstrations; metrics show the mean over five seeds with 95% confidence intervals.
@@ 6 Conclusion
We propose a novel IRL framework that overcomes the limitations of fixed reward mechanisms through dynamic reward adaptation and context-sensitive regularization. Our approach ensures bounded implicit rewards and stable value function updates, leading to robust policy optimization. By integrating distributional RL with implicit reward learning, we capture richer return dynamics while preserving theoretical guarantees. Empirical results on MuJoCo and Adroit benchmarks show expert-like proficiency on the Humanoid-v2 and Hammer-v1 tasks with three expert demonstrations. Ablation studies confirm the important role of our regularization mechanism. This work unifies implicit reward regularization with distributional return representations, offering a scalable and sample-efficient solution for complex decision-making. Future directions include extending this framework to offline IL and risk-sensitive robotics, where TD regularizer with adaptive targets and uncertainty-aware return distributions can further improve robustness and generalization.
@@ Appendix A Supporting Proofs
A.1 Expectation of the Return Distribution
@@
From the soft return distribution definition
@@
A.2 Convergence of the Optimal Reward
@@
Appendix B Implementation Details
@@
MuJoCo Suite. We evaluate RIZE on five Gym (Brockman2016) MuJoCo (Todorov2012) locomotion tasks: HalfCheetah-v2, Walker2d-v2, Ant-v2, Humanoid-v2, and Hopper-v2. Expert trajectories for these tasks are taken from IQ-Learn (Garg2021) and were generated with Soft Actor–Critic (Haarnoja2018); each trajectory contains 1,000 state–action transitions. Episode returns are normalized by expert performance with the following expert evaluation returns: HalfCheetah (5,100), Walker2d (5,200), Ant (4,700), Humanoid (5,300), and Hopper (3,500).
Adroit. For Hammer-v1 from the Adroit suite (rajeswaran2018dexterousmanipulation), we use the D4RL dataset (fu2021d4rldatasetsdeepdatadriven) and filter the top 100 episodes from the original 5,000. The resulting expert subset has an average return of 16,800, and each episode comprises 200 state–action pairs. We normalize returns in this domain by the average expert return of the selected subset.
Baselines. We evaluate five baselines: IQ-Learn (Garg2021), LSIQ (Al-Hafez2023), SQIL (Reddy2020), CSIL (Watson2023), and Behavior Cloning (BC) (Pomerleau1991). For MuJoCo tasks, we use the authors’ original code and configurations for IQ-Learn, LSIQ, and CSIL; SQIL is implemented using the LSIQ codebase, and BC follows the CSIL implementation. For Hammer-v1, we use the original CSIL code and configs. For IQ-Learn on Hammer, we observe that it does not solve the task even after a small sweep over the entropy coefficient in $\{0.01,0.03,0.1\}$ and the loss mode in {value, v0}; we report $\alpha{=}0.03$ and loss=value. For LSIQ on Hammer, we search over $\alpha\in\{0.01,0.05,0.1\}$ and loss $\in\{\texttt{value},\texttt{v0}\}$ and select $\alpha{=}0.1$, loss=value as the better-performing variant. For SQIL, we set $\alpha{=}0.2$ consistent with our other tasks.
@@
Our architecture integrates components from Distributional SAC (DSAC) [^1] https://github.com/xtma/dsac (Ma2020) and IQ-Learn [^2] https://github.com/Div99/IQ-Learn (Garg2021), with hyperparameters tuned through search and ablation studies. Key configurations for experiments involving three and ten demonstrations are summarized in Table 1. All implementation details used in our experiments are publicly available at https://github.com/adibka/RIZE.
@@
Distributional SAC Components. The critic network is implemented as a three-layer multilayer perceptron (MLP) with 256 units per layer, trained using a learning rate of $3\times 10^{-4}$. The policy network is a four-layer MLP, also with 256 units per layer. To enhance training stability, we employ a target policy—a delayed version of the online policy—and sample next-state actions from this module. For return distribution training $Z^{\pi}_{\phi,\tau}$, we adopt the Implicit Quantile Networks (IQN) (Dabney2018) approach by sampling quantile fractions $\tau$ uniformly from $\mathcal{U}(0,1)$. Additionally, dual critic networks with delayed updates are used, which empirically improve training stability. We use the following settings across tasks: replay buffer size $10^{6}$, batch size 256, 24 quantile levels, and 10,00010{,}000 pretraining steps. We evaluate every $10^{4}$ steps, which takes  3.5 minutes.
@@
IQ-Learn Adaptations. Key adaptations from IQ-Learn include adjustments to the regularizer coefficient $c$ and entropy coefficient $\alpha$. Specifically, for the regularizer coefficient $c$, we find that $c=0.5$ yields robust performance on the Humanoid task, while $c=0.1$ works better for other tasks. For the entropy coefficient $\alpha$, smaller values lead to more stable training. Unlike RL, where exploration is crucial, imitation learning relies less on entropy due to the availability of expert data. Across all tasks, we set initial target reward parameters as $\lambda^{\pi_{E}}=10$ and $\lambda^{\pi}=5$. Furthermore, we observe that lower learning rates for target rewards improve overall learning performance.
@@
The choice between v0 or value loss variants depends on environment complexity: we find that for a complex task like Humanoid-v2, the v0 variant demonstrates greater robustness. And, for HalfCheetah-v2, Walker2d-v2, Hopper-v2, Ant-v2, and Hammer-v1, the value variant performs better.